OpenStudy (anonymous):

find x and y using substitution x=6y-4 x=-2y+28 what to do next?

OpenStudy (amistre64):

This is pretty simple really: since both equation = x, we can set them equal to each other.

OpenStudy (amistre64):

6y -4 = -2y +28 get like terms together: 6y +2y = 28 +4 combine the like terms: 8y = 32 get y all by its little lonesome by dividing both sides by (8): 8y/8 = 32/8 y = 4

OpenStudy (amistre64):

double check to make sure and: x=6(4)-4 x = 24 - 4 = 20 x=-2(4)+28 = -8 + 28 = 20 were good to go....